melissaa brother is six years less than twice her age. the sum of her age and her brother's age is twenty seven. what should the equation be?
Answer by sunbern(1)

M(Melissa's Age)
B(Brother's Age)
B=2M - 6
M + B = 27
Melissa's Age = 11
Melissa's Brother's Age = 16
16 = 2(11) - 6
11 + 6 = 7